# Boolean Nearest Neighbor Language in the Knowledge Compilation Map

### v1

@@ Abstract
The Boolean Nearest Neighbor (BNN) representation of Boolean functions was recently introduced by Hajnal, Liu and Turán [13]. A BNN representation of $f$ is a pair $(P,N)$ of sets of Boolean vectors (called positive and negative prototypes) where $f(x)=1$ for every positive prototype $x\in P$, $f(x)=0$ for all every negative prototype $x\in N$, and the value $f(x)$ for $x\not\in P\cup N$ is determined by the type of the closest prototype. The main aim of this paper is to determine the position of the BNN language in the Knowledge Compilation Map (KCM). To this end, we derive results which compare the succinctness of the BNN language to several standard languages from KCM, and determine the complexity status of most standard queries and transformations for BNN inputs.
@@ 1 Introduction
The task of transforming one of the representations of a given function $f$ into another representation of $f$ (e.g. transforming a DNF representation into an OBDD or a DNNF into a CNF) is called knowledge compilation. This task emerged as an important ingredient of modern computation, aiming to transform Boolean functions into more compact and tractable forms. This allows us to use techniques from logic, algorithms and data structures in order to bridge the gap between the expressive power of Boolean functions and the efficiency requirements of practical applications. For a comprehensive review paper on knowledge compilation see [10], where the Knowledge Compilation Map (KCM) is introduced. KCM systematically investigates different knowledge representation languages with respect to (1) their relative succinctness, (2) the complexity of common transformations, and (3) the complexity of common queries. The succinctness of representations roughly speaking describes how large the output representation in language $B$ is with respect to the size of the input representation in language $A$ when compiling from $A$ to $B$. A precise definition of this notion will be given later in this text. Transformations include negation, conjunction, disjunction, conditioning, and forgetting. The complexity of such transformations may differ dramatically from trivial to computationally hard depending on the chosen representation language. The same is true for queries such as consistency check, validity check, clausal and sentential entailment, equivalence check, model counting, and model enumeration.
@@
BNN representations of Boolean functions were introduced in [13] as a special case of a more general Nearest Neighbor (NN) representations where the prototypes are not restricted to Boolean vectors but instead can be any real-valued vectors from $\mathbb{R}^{n}$. In this case Eucleidian distance $d_{E}$ is used instead of the Hamming distance $d_{H}$. Notice that there is in fact no difference between using Euclidean and Hamming distance when considering Boolean prototypes. This is due to the fact that that for any two vectors $x,y\in\mathcal{B}^{n}$, $d_{E}(x,y)=\sqrt{d_{H}(x,y)}$. Thus any BNN representation of a function $f$, is also an NN representation of $f$. NN representations of Boolean functions were already studied more than thirty years ago [19], but there are also many recent works [2, 12, 14].
@@ 2 Definitions and Recent Results
Another easy consequence of the previous two lemmas concerns isolated vectors. Vector $x\in\mathcal{B}^{n}$ is called isolated for $f$ if all its neighbors have the opposite function value (i.e. $x$ represents a connected components of size one in the corresponding subgraph).
@@
Let $f$ be a Boolean function and let $x$ be positive (resp. negative) vector such that $x$ is isolated for $f$. Then $x$ must be a positive (resp. negative) prototype in any BNN representation of $f$.
@@ 3 Succinctness
Let $L_{1}$ and $L_{2}$ be two knowledge representations languages. We say that $L_{1}$ is at least as succinct as $L_{2}$, denoted $L_{1}\leq L_{2}$, if and only if there exists a polynomial $p$ such that for every sentence $\alpha\in L_{2}$, there exists an equivalent sentence $\beta\in L_{1}$ where $|\beta|\leq p(|\alpha|)$. Furthermore, $L_{1}$ is strictly more succinct than $L_{2}$, denoted $L_{1}<L_{2}$, if $L_{1}\leq L_{2}$ but $L_{2}\not\leq L_{1}$.
@@
For the non-strict relation BDD $\leq$ BNN it suffices to show that there exists a polynomial $p$ such that for every sentence $\alpha\in\mathbf{BNN}$, there exists an equivalent sentence $\beta\in\mathbf{BDD}$ where $|\beta|\leq p(|\alpha|)$. For any $\alpha=(P,N)$, we will construct such a a binary decision diagram $\beta$. Let us assume that $P\cup N=\{p^{1},\dots,p^{k}\}$ and let $x\in\mathcal{B}^{n}$. We build $\beta$ in two steps:
@@
Consider for an example the gadget in Figure 3 with $x=(1,0,1)$ and $p^{i}=(0,0,1)$. In the first three coordinated comparisons, we take the edges labelled $N$, $Y$ and $Y$. As only one coordinate differs, on level $4$ we know that the value of $d_{H}(x,p^{i})$ is $1$.
@@
It now remains to build the output BDD $\beta$ using the gadgets. We can sequentially test pairs of prototypes until the closest one is found. We do so by making a triangle shaped acyclic directed graph made of gadgets with $k-1$ levels. At level $i$, prototype $p^{i+1}$ is compared with every $p^{j}$ for $j\leq i$, and directed edges are pointed to appropriate gadgets on the next level. After level $k-1$, a closest prototype has been determined, and so the directed edges going out of these gadgets point to terminal $1$ (resp. 00) depending on whether the found closest prototype is positive (resp. negative). The output BDD $\beta$ for a function with $k$ prototypes is shown in Figure 5.
@@
Consider the majority function $g_{n}$ on $n$ variables which can be represented by a BNN of size $O(n)$ by Theorem 2.7. On the other hand any CNF of this monotone function must contain at least as many clauses as is the number of maximal false vectors of $g_{n}$ [8] which is of course exponential in $n$, and so the claim follows. ∎
@@
It suffices to show that there is a Boolean function in $n$ variables with OBDD representation of polynomial size in $n$ which cannot be represented by a BNN of polynomial size in $n$. This property is satisfied by the threshold function $TH_{n}^{\lfloor n/3\rfloor}$ which is true if and only if at least one third of its inputs is set to one. This is a symmetric Boolean function and it is long known that every such function can be represented by an OBDD of polynomial size in the number of variables [18]. On the other hand, $TH_{n}^{\lfloor n/3\rfloor}$ can only be represented by a BNN of exponential size in $n$ by Theorem 2.7. ∎
@@ 4 Transformations
Let $(P,N)$ be the smallest $BNN$ representing the Boolean majority function on $n=4k$ variables, for some $k\in\mathbb{N}$. That is, $(P,N)$ represents the function $TH_{4k}^{2k}$ and $|(P,N)|\leq\frac{n}{2}+2=2(k+1)$ holds by Theorem 2.7. Let $x_{1},\dots,x_{n}$ denote the variables of the threshold function. Notice that by setting some variable $x_{i}$ to $1$, i.e. conditioning on term $T=x_{i}$ we obtain a threshold function which has one variable fewer and threshold smaller by one.
@@
This function has a BNN representation of size $O(n)$ by Theorem 2.7. Furthermore consider the (non-strict) minority function on $2n$ variables
@@ 5 Queries
In this section we shall show that the BNN language supports a reasonably large subset of standard queries from [10]. We begin with a trivial observation that both consistency and validity can be checked in constant time.
@@
Next we prove that implicant check is supported by the BNN language.
@@
Let $f$ be a Boolean function and $(P,N)$ its BNN representation. Let $T=l_{1}\land\dots\land l_{k}$ be a consistent term. Without loss of generality, we may assume that $\forall 1\leq i\leq k:$ $l_{i}\in\{x_{i},\lnot x_{i}\}$, since otherwise we may relabel the variables. Our aim is to design a polynomial time algorithm which checks whether $T\implies f$. Let us denote by $y\in\mathcal{B}^{k}$ the (partial) vector satisfying $T$, i.e. defined by $y_{i}=1$ if $l_{i}=x_{i}$ and $y_{i}=0$ if $l_{i}=\lnot x_{i}$. Furthermore, let $S$ denote the sub-cube of $\mathcal{B}^{n}$ determined by the vector $y$. That is,
@@
The above discussion is summarized in Algorithm 1. The algorithm runs in time $O(n|P||N|)$ (which is polynomial in the size of the input $n(|P|+|N|)$), as its main part consists of two nested for loops. In the worst case, the algorithms considers all projections of negative prototypes, and calculates their distances to each of the positive prototypes. ∎
@@
Let $f$ be a Boolean function and $(P,N)$ its BNN representation. Let $C=l_{1}\lor\dots\lor l_{k}$ be a consistent clause. Our aim is to design a polynomial time algorithm which checks whether $f\implies C$. Clearly $(f\implies C)\equiv(\lnot C\implies\lnot f)$. Moreover, $\lnot f$ is readily available (recall Observation 4.1), and by DeMorgan laws the negation of a clause $C$ is a term $T$. So let $T=\lnot C$. Now, we may check whether $f$ entails $C$ by calling Algorithm 1 for inputs $\lnot f=(N,P)$ and $\lnot C=T$, getting the correct answer in polynomial time. ∎
@@
Models can be enumerated using a tree search such as the simple DPLL algorithm where at every step before a value is assigned to a variable and a branch to a node on the next level is build, it is first tested whether the current partial assignment of values to variables plus the considered assignment yields a satisfiable formula (and if not then the branch is not build). This of course amounts to a clausal entailment test. Therefore all branches of the tree that is built have full length (all variables are fixed to constants) and terminate at models. Thus both the size of the tree and the total work required is upper bounded by a polynomial in the number of models. ∎
A question whether BNN supports any of the other queries remains open. We conjecture the remaining standard queries are hard, i.e. that there are no polynomial time algorithms for EQ, SE and CT (unless P=NP). The complexity status of all standard queries for BNN and selected standard languages is summarized in Table 2.

### v2

@@ Abstract
The Boolean Nearest Neighbor (BNN) representation of Boolean functions was recently introduced by Hajnal, Liu and Turán [13]. A BNN representation of $f$ is a pair $(P,N)$ of sets of Boolean vectors (called positive and negative prototypes) where $f(x)=1$ for every positive prototype $x\in P$, $f(x)=0$ for all every negative prototype $x\in N$, and the value $f(x)$ for $x\not\in P\cup N$ is determined by the type of the closest prototype. The main aim of this paper is to determine the position of the BNN language in the Knowledge Compilation Map (KCM). To this end, we derive results that compare the succinctness of the BNN language to several standard languages from KCM and determine the complexity status of most standard queries and transformations for BNN inputs.
@@ 1 Introduction
The task of transforming one of the representations of a given function $f$ into another representation of $f$ (e.g. transforming a DNF representation into an OBDD or a DNNF into a CNF) is called knowledge compilation. This task emerged as an important ingredient of modern computation, aiming to transform Boolean functions into more compact and tractable forms. This allows us to use techniques from logic, algorithms, and data structures in order to bridge the gap between the expressive power of Boolean functions and the efficiency requirements of practical applications. For a comprehensive review paper on knowledge compilation see [10], where the Knowledge Compilation Map (KCM) is introduced. KCM systematically investigates different knowledge representation languages with respect to (1) their relative succinctness, (2) the complexity of common transformations, and (3) the complexity of common queries. The succinctness of representations roughly speaking describes how large the output representation in language $B$ is with respect to the size of the input representation in language $A$ when compiling from $A$ to $B$. A precise definition of this notion will be given later in this text. Transformations include negation, conjunction, disjunction, conditioning, and forgetting. The complexity of such transformations may differ dramatically from trivial to computationally hard depending on the chosen representation language. The same is true for queries such as consistency check, validity check, clausal and sentential entailment, equivalence check, model counting, and model enumeration.
@@
BNN representations of Boolean functions were introduced in [13] as a special case of a more general Nearest Neighbor (NN) representations where the prototypes are not restricted to Boolean vectors but instead can be any real-valued vectors from $\mathbb{R}^{n}$. In this case, Euclidean distance $d_{E}$ is used instead of the Hamming distance $d_{H}$. Notice that there is in fact no difference between using Euclidean and Hamming distance when considering Boolean prototypes. This is due to the fact that for any two vectors $x,y\in\mathcal{B}^{n}$, $d_{E}(x,y)=\sqrt{d_{H}(x,y)}$. Thus any BNN representation of a function $f$, is also an NN representation of $f$. NN representations of Boolean functions were already studied more than thirty years ago [19], but there are also many recent works [2, 12, 14].
@@ 2 Definitions and Recent Results
Another easy consequence of the previous two lemmas concerns isolated vectors. Vector $x\in\mathcal{B}^{n}$ is called isolated for $f$ if all its neighbors have the opposite function value (i.e. $x$ represents a connected component of size one in the corresponding subgraph).
@@
Let $f$ be a Boolean function and let $x$ be a positive (resp. negative) vector such that $x$ is isolated for $f$. Then $x$ must be a positive (resp. negative) prototype in any BNN representation of $f$.
@@ 3 Succinctness
Let $L_{1}$ and $L_{2}$ be two knowledge representation languages. We say that $L_{1}$ is at least as succinct as $L_{2}$, denoted $L_{1}\leq L_{2}$, if and only if there exists a polynomial $p$ such that for every sentence $\alpha\in L_{2}$, there exists an equivalent sentence $\beta\in L_{1}$ where $|\beta|\leq p(|\alpha|)$. Furthermore, $L_{1}$ is strictly more succinct than $L_{2}$, denoted $L_{1}<L_{2}$, if $L_{1}\leq L_{2}$ but $L_{2}\not\leq L_{1}$.
@@
For the non-strict relation BDD $\leq$ BNN it suffices to show that there exists a polynomial $p$ such that for every sentence $\alpha\in\mathbf{BNN}$, there exists an equivalent sentence $\beta\in\mathbf{BDD}$ where $|\beta|\leq p(|\alpha|)$. For any $\alpha=(P,N)$, we will construct such a binary decision diagram $\beta$. Let us assume that $P\cup N=\{p^{1},\dots,p^{k}\}$ and let $x\in\mathcal{B}^{n}$. We build $\beta$ in two steps:
@@
Consider for an example the gadget in Figure 3 with $x=(1,0,1)$ and $p^{i}=(0,0,1)$. In the first three coordinated comparisons, we take the edges labeled $N$, $Y$ and $Y$. As only one coordinate differs, on level $4$ we know that the value of $d_{H}(x,p^{i})$ is $1$.
@@
It now remains to build the output BDD $\beta$ using the gadgets. We can sequentially test pairs of prototypes until the closest one is found. We do so by making a triangle-shaped acyclic directed graph made of gadgets with $k-1$ levels. At level $i$, prototype $p^{i+1}$ is compared with every $p^{j}$ for $j\leq i$, and directed edges are pointed to appropriate gadgets on the next level. After level $k-1$, a closest prototype has been determined, and so the directed edges going out of these gadgets point to terminal $1$ (resp. 00) depending on whether the found closest prototype is positive (resp. negative). The output BDD $\beta$ for a function with $k$ prototypes is shown in Figure 5.
@@
Consider the majority function $g_{n}$ on $n$ variables which can be represented by a BNN of size $O(n)$ by Theorem 2.7. On the other hand, any CNF of this monotone function must contain at least as many clauses as is the number of maximal false vectors of $g_{n}$ [8] which is of course exponential in $n$, and so the claim follows. ∎
@@
It suffices to show that there is a Boolean function in $n$ variables with OBDD representation of polynomial size in $n$ which cannot be represented by a BNN of polynomial size in $n$. This property is satisfied by the threshold function $TH_{n}^{\lfloor n/3\rfloor}$ which is true if and only if at least one third of its inputs are set to one. This is a symmetric Boolean function and it is long known that every such function can be represented by an OBDD of polynomial size in the number of variables [18]. On the other hand, $TH_{n}^{\lfloor n/3\rfloor}$ can only be represented by a BNN of exponential size in $n$ by Theorem 2.7. ∎
@@ 4 Transformations
Let $(P,N)$ be the smallest $BNN$ representing the Boolean majority function on $n=4k$ variables, for some $k\in\mathbb{N}$. That is, $(P,N)$ represents the function $TH_{4k}^{2k}$ and $|(P,N)|\leq\frac{n}{2}+2=2(k+1)$ holds by Theorem 2.7. Let $x_{1},\dots,x_{n}$ denote the variables of the threshold function. Notice that by setting some variable $x_{i}$ to $1$, i.e. conditioning on term $T=x_{i}$ we obtain a threshold function that has one variable fewer and threshold smaller by one.
@@
This function has a BNN representation of size $O(n)$ by Theorem 2.7. Furthermore, consider the (non-strict) minority function on $2n$ variables
@@ 5 Queries
In this section, we shall show that the BNN language supports a reasonably large subset of standard queries from [10]. We begin with a trivial observation that both consistency and validity can be checked in constant time.
@@
Next, we prove that implicant check is supported by the BNN language.
@@
Let $f$ be a Boolean function and $(P,N)$ its BNN representation. Let $T=l_{1}\land\dots\land l_{k}$ be a consistent term. Without loss of generality, we may assume that $\forall 1\leq i\leq k:$ $l_{i}\in\{x_{i},\lnot x_{i}\}$, since otherwise we may relabel the variables. Our aim is to design a polynomial time algorithm that checks whether $T\implies f$. Let us denote by $y\in\mathcal{B}^{k}$ the (partial) vector satisfying $T$, i.e. defined by $y_{i}=1$ if $l_{i}=x_{i}$ and $y_{i}=0$ if $l_{i}=\lnot x_{i}$. Furthermore, let $S$ denote the sub-cube of $\mathcal{B}^{n}$ determined by the vector $y$. That is,
@@
The above discussion is summarized in Algorithm 1. The algorithm runs in time $O(n|P||N|)$ (which is polynomial in the size of the input $n(|P|+|N|)$), as its main part consists of two nested for loops. In the worst case, the algorithm considers all projections of negative prototypes and calculates their distances to each of the positive prototypes. ∎
@@
Let $f$ be a Boolean function and $(P,N)$ its BNN representation. Let $C=l_{1}\lor\dots\lor l_{k}$ be a consistent clause. Our aim is to design a polynomial time algorithm that checks whether $f\implies C$. Clearly $(f\implies C)\equiv(\lnot C\implies\lnot f)$. Moreover, $\lnot f$ is readily available (recall Observation 4.1), and by DeMorgan laws the negation of a clause $C$ is a term $T$. So let $T=\lnot C$. Now, we may check whether $f$ entails $C$ by calling Algorithm 1 for inputs $\lnot f=(N,P)$ and $\lnot C=T$, getting the correct answer in polynomial time. ∎
@@
Models can be enumerated using a tree search such as the simple DPLL algorithm where at every step before a value is assigned to a variable and a branch to a node on the next level is built, it is first tested whether the current partial assignment of values to variables plus the considered assignment yields a subfunction which is identically zero (and if yes then the branch is not built). This of course amounts to a clausal entailment test. Therefore all branches of the tree that is built have full length (all variables are fixed to constants) and terminate at models. Thus both the size of the tree and the total work required is upper bounded by a polynomial in the number of models. ∎
In the rest of this section, we shall prove that BNN does not support the remaining standard queries, i.e. that there are no polynomial time algorithms for EQ, SE and CT queries unless P=NP. All three proofs are based on the same reduction from the following NP-complete decision problem:
Half-Size Independent Set (HSIS)
Input: An undirected graph $G=(V,E)$ with $n$ vertices.
Question: Does there exist an independent set with exactly $n/2$ vertices in $G$?
Although IS, the general independent set problem (in which a parameter $k$ is part of the input and the question asks for the existence of an independent set of size $k$), is widely known to be NP-complete, we have to argue that also the restricted version HSIS is hard. To see this, consider the textbook reduction from 3-SAT to IS which for every cubic clause creates a clique of size three and then connects vertices that correspond to complementary literals (one edge for each such pair). It is easy to see that the input 3-SAT instance (with $m$ clauses) is satisfiable if and only if the constructed graph on $3m$ vertices contains an independent set of size exactly $m$. Modifying this construction by adding $m$ isolated vertices yields a reduction in which the input 3-SAT instance is satisfiable if and only if the constructed graph on $4m$ vertices contains an independent set of size exactly $2m$, which is an instance of HSIS.
Now we are ready to prove the first hardness result.
Theorem 5.5.
The equivalence query is co-NP complete for the BNN language.
Proof.
It is easy to see that the EQ query for BNN belongs to co-NP. A certificate for a negative answer is a vector on which the two input BNN representations give opposite function values (which can be of course checked in polynomial time).
For the hardness part, let $G=(V,E)$ with $n=2k$ vertices be an instance of HSIS (we assume without loss of generality that $k>2$). We define two input BNN representations for the EQ query as follows:
  1. 1.
    $F=(P_{f},N_{f})$ is a BNN representation of the majority function $f=MAJ_{2k}$ on $2k$ variables. We assume here that $F$ is the representation of $f$ from the proof of Theorem 2.7, namely
    • $P_{f}=\{x\in\mathcal{B}^{2k}\mid|x|=2k-1\}$, i.e all [^1] The proof of Theorem 2.7 in fact uses only arbitrarily chosen $k+1$ vectors of weight $12k-$ 1, however, we do not need a minimal representation here and taking all vectors of weight $12k-$ 1 does not change the represented function. vectors of weight $12k-$ 1, and
    • $N_{f}=\{(0,0,\ldots,0)\}$, i.e. the only vector of weight 00.
  2. 2.
    $H=(P_{h},N_{h})$ is a BNN representation of function $h$ defined as follows
    • $P_{h}=\{p\}\cup\{p^{e}\mid e\in E\}$, where $p=(1,1,\ldots,1)$ is the only vector of weight $2k$ and for every $e=(i,j)\in E$ the vector $p^{e}$ has weight $22k-$ 2 with $p^{e}_{i}=p^{e}_{j}=0$ (all remaining $22k-$ 2 bits in $p^{e}$ are $1$).
    • $N_{h}=\{x\in\mathcal{B}^{2k}\mid|x|=1\}$, i.e all vectors of weight $1$. Let us denote these vectors by $n^{1},\ldots,n^{2k}$ with $n^{i}_{i}=1$ for every $1\leq i\leq 2k$ (all remaining $12k-$ 1 bits in $n^{i}$ are 00).
First, we shall show that $h(x)=f(x)$ holds for all vectors with weight different from $k$.
  • Let $x\in\mathcal{B}^{2k}$ be an arbitrary vector with $|x|\leq k-1$. Clearly, if we pick any index $i$ with $x_{i}=1$ then the negative prototype $n^{i}$ of weight $1$ satisfies $d_{H}(x,n^{i})\leq k-2$ (and if $x$ is the all-zero vector then $d_{H}(x,n^{i})=1\leq k-2$ for every $i$ because $k>2$ was assumed). On the other hand, all positive prototypes are at a Hamming distance at least $k-1$ from $x$ since at least that many zero-bits in $x$ have to be flipped to get a vector of weight $22k-$ 2 or more. Thus $h(x)=0=f(x)$.
  • Let $x\in\mathcal{B}^{2k}$ be an arbitrary vector with $|x|\geq k+1$. In this case, we have $d_{H}(x,p)\leq k-1$ while all negative prototypes are at a Hamming distance at least $k$ from $x$ since at least that many one-bits in $x$ have to be flipped to get a vector of weight $1$. Thus $h(x)=1=f(x)$.
Thus $h(x)=f(x)$ holds everywhere except of the middle level $|x|=k$ of the Boolean lattice which we shall investigate now.
  • Assume that there exists no independent set of size $k$ in $G$ and let $x\in\mathcal{B}^{2k}$ be an arbitrary vector with $|x|=k$. Vector $x$ has exactly $k$ coordinates with $x_{i}=0$ and by our assumption, this index set cannot be an independent set of $G$. Thus there must exist $e=(i,j)\in E$ such that $x_{i}=x_{j}=0$ and hence $d_{H}(x,p^{e})=k-2$ since only the remaining $k-2$ zero-bits in $x$ have to be flipped to arrive to $p^{e}$. On the other hand, all negative prototypes are at a Hamming distance at least $k-1$ from $x$ since at least that many one-bits in $x$ have to be flipped to get a vector of weight $1$. Thus $h(x)=1=f(x)$ for every vector $x$ of weight $k$.
  • Now assume the opposite, i.e. that there exists an independent set of size $k$ in $G$, defined by an index set $S$. Let $x^{S}\in\mathcal{B}^{2k}$ be a vector with $|x^{S}|=k$ where $S$ defines the zero-bits of $x^{S}$. Clearly, $d_{H}(x^{S},p)=k$ but also $d_{H}(x^{S},p^{e})\geq k$ for every $e\in E$. To see the latter, notice that if $|e\cap S|=1$ we need to flip $k-1$ zero-bits and one one-bit in $x^{S}$ to arrive to $p^{e}$, and if $|e\cap S|=0$ we need to flip all $k$ zero-bits and two one-bits in $x^{S}$ to arrive to $p^{e}$. On the other hand, if we pick any index $i$ with $x^{S}_{i}=1$ then the negative prototype $n^{i}$ of weight $1$ satisfies $d_{H}(x^{S},n^{i})=k-1$. Thus $h(x^{S})=0$ while $f(x^{S})=1$.
If we summarize the above observations we get that $h\equiv f$ if and only if there exists no independent set of size $k$ in $G$, or in other words the answer to the EQ query on BNN representations $F$ and $H$ is yes if and only if the answer to the input HSIS instance $G$ is no. Since both $F$ and $H$ can be clearly constructed from $G$ in polynomial time, this finishes the proof of NP-harness of the EQ query. ∎
Theorem 5.5 immediately gives us the following corollary.
Corollary 5.6.
The sentential entailment query is co-NP complete for the BNN language.
Proof.
It is again easy to see that the SE query for BNN belongs to co-NP. Given a query $F\models H$ a certificate for a negative answer is a vector $x$ such that $F$ classifies $x$ as a positive vector while $H$ classifies $x$ as a negative vector (which can be of course checked in polynomial time).
The hardness part is a direct consequence of Theorem 5.5 since $F\equiv H$ if and only if $F\models H$ and $H\models F$ so any EQ query can be answered by asking two SE queries on the same input BNN representations. ∎
Finally, the construction in the proof of Theorem 5.5 also yields the following result.
Corollary 5.7.
The model counting query is NP-hard for the BNN language.
Proof.
[…]
It is also obvious, that $CT(H)=CT(F)$ if and only if $H\equiv F$ (recall that the models of $H$ always form a subset of models of $F$) and so the ability to compute $CT(H)$ in polynomial time would answer the EQ query for $F$ and $H$ and thus decide the input HSIS instance $G$. ∎
Let us remark that the reduction in the proof of Theorem 5.5 is number preserving in the following sense: any two distinct independent sets of size $k$ in $G$ correspond to two distinct non-models of $h$ of weight $k$. Therefore $CT(F)-CT(H)$ exactly equals the number of independent sets of size $k$ in $G$. This means that if the counting version of the HSIS problem is #P hard (which we do not know, but it is quite likely since the counting version of the general IS problem certainly is #P hard) then also the CT query for BNN is #P hard.
The complexity status of all standard queries for BNN and selected standard languages is summarized in Table 2.